# New analysis of SNeIa Pantheon Catalog:  Variable speed of light as   an alternative to dark energy

### v2

@@ II A new mechanism to generate varying c from dilaton dynamics
Einstein’s pioneering concept of VSL, nevertheless, was quickly overshadowed by the success of his GR and subsequently fell into dormancy for several decades. The variability of $c$ was briefly rediscovered by Dicke in 1957 [11], prior to his own development of Brans–Dicke gravity [12], which instead allowed Newton’s gravitational constant $G$ to vary. In the 1990s, the idea of VSL was independently revived by Moffat [13] and by Albrecht and Magueijo [14] in the context of early-time cosmology. Their proposals aimed to resolve the horizon puzzle while avoiding the need for cosmic inflation. Since then, several researchers actively explore various aspects of VSL [33, 34, 35, 17, 27, 30, 29, 28, 18, 19, 49, 31, 32, 47, 48, 65, 80, 68, 45, 44, 43, 63, 62, 37, 36, 58, 72, 70, 75, 51, 71, 46, 25, 76, 81, 77, 73, 78, 55, 64, 66, 69, 39, 38, 52, 74, 79, 40, 53, 50, 54, 82, 26, 24, 22, 67, 61, 60, 21, 20, 16, 15, 57, 56, 23, 42, 41, 59, 83, 84, 85, 86].
@@ III.2 The modified FLRW metric
Here, $a_{0}$ is the current cosmic scale factor (often set equal 1), and $c_{0}$ is the speed of light measured at our current time in the intergalactic space. We should emphasize that the value of $c_{0}$ is not identical with the one measured inside the Milky Way, which is equal to $300,000\ \text{km sec}^{-1}$. This is because the Milky Way is a gravitationally bound structure whereas the intergalactic space is regions subject to cosmic expansion. This issue will be explained in Section IV.
@@ V.4 Modifying the luminosity distance–redshift formula
$4\pi r^{2}J=L\frac{\hat{\lambda}_{ob}}{\hat{\lambda}_{em}}.\frac{\hat{\lambda}_{ob}}{\hat{\lambda}_{em}}$ (47)
In the right hand side of Eq. (47), the first term $\hat{\lambda}_{ob}/\hat{\lambda}_{em}$ represents the “loss” in energy of the redshifted photon known as the “Doppler theft”  [^2] Note: This energy loss is consistent with the scaling property of energy for our VSL mechanism, as described by Eq. (13) in Section II. In intergalactic space, the energy of the traveling photon scales as $E\propto\chi\propto a^{-1}$, leading to a decline in energy as the universe expands.. The second (identical) term $\hat{\lambda}_{ob}/\hat{\lambda}_{em}$ arises from the dilution factor in photon density, as the same number of photons is distributed over a prolonged wavecrest in the radial direction (i.e., along the light ray). The $4\pi r^{2}$ in the left hand side of Eq. (47) accounts for the spherical dilution in flat space. From (46) and (47), we obtain
$d_{L}=r\,\frac{\hat{\lambda}_{ob}}{\hat{\lambda}_{ob}}$ (48)
@@
where $\hat{d}_{L}$ is the luminosity distance observed by the Earth-based astronomer and $\hat{c}_{ob}$ the speed of light measured in the Milky Way (i.e., $300,000$ km sec-1). Formula (52) contains a single parameters $H_{0}$ and involves a function $F(z)$ that captures the evolution of the local scale of galaxies as a function of redshift.
@@ VI.1 $\Lambda$CDM and standard EdS as benchmarking models
with the sum taken over all $N=1,048$ Pantheon data points. The best fit for the $\Lambda$CDM model yields $H_{0}=70.2$ km sec-1 Mpc-1, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$, with the minimum error $\chi_{\text{min}}^{2}(\Lambda\text{CDM})=0.98824$. The $d_{L}$$z$ curve for the $\Lambda$CDM model is depicted by the dashed line in Fig. 3.
@@ VI.2 Fitting with VSL model: Disabling $F(z)$
The best fit of the Pantheon data to this formula yields $H_{0}=44.4\text{ km s}^{-1}\text{ Mpc}^{-1}$, corresponding to $\chi_{\text{min}}^{2}=1.25366$. Figure 3 displays our fit as the solid line. Although this fit performs worse than the $\Lambda$CDM model, which has $\chi_{\text{min}}^{2}(\Lambda\text{CDM})=0.98824$, it substantially reduces the excess in distance moduli for $z\gtrsim 0.1$ compared with the “fiducial” EdS model, as shown in Fig. 3.
@@ VI.3 Enabling $F(z)$:   A binning approach
Instead of fitting each bin separately, we impose one common value for $H_{0}$ across all bins. The fit thus involves $H_{0}$ and 10 values for $\{F^{(i)},\,i=1..10\}$. Figure 4 displays our fit to Formula (60). The best fit yields $H_{0}=46.6\text{ km s}^{-1}\text{ Mpc}^{-1}$, and the values of $F^{(i)}$ are given in the last column of Table 1. The minimum error is $\chi_{\text{min}}^{2}=0.97803$. The staircase lines in Fig. 5 depict $F^{(i)}$ as function of $z$ and the cosmic factor $a$.
@@
VI.4 Enabling $F(z)$:   A functional form approach
@@
The steady decline of $F^{(i)}$ across the 10 bins with respect to redshift suggests adopting the following functional form for $F(z)$:
@@
Figure 5 shows the variation of $F$ in dashed lines with respect to both redshift and the cosmic factor. We also produce the joint distribution for $H_{0}$ and $F_{\infty}$, as shown in Figure 7, yielding $H_{0}=47.21\pm 0.4\text{ km s}^{-1}\text{ Mpc}^{-1}$ (95% CL) and $F_{\infty}=0.931\pm 0.008$ (95% CL). This value of $F_{\infty}$ indicates that the local scales of galaxies have increased by approximately 7% since the formation of the first stable galaxies (i.e., those at the largest redshift).
@@ VI.5 The cause for the reduction in $H_{0}$ value
Using this new function, BDRS produced an excellent fit to the CMB power spectrum, resulting in the following parameters: $H_{0}=46$ km sec-1 Mpc-1, $\omega_{\text{baryon}}:=\Omega_{\text{baryon}}(H_{0}/100)^{2}=0.019$, $\tau=0.16$ (the optical depth to last scattering), $k_{*}=0.0096$ Mpc-1, $n_{1}=1.015$, and $n_{2}=0.806$. The most remarkable outcome of the BDRS work is the “low” value of $H_{0}=46$, representing a 34% reduction from the accepted value of $H_{0}\sim 70$. A detailed follow-up study by Hunt and Sarkar [5], based on a supergravity-induced multiple inflation scenario, yielded a comparable value of $H_{0}\approx 44$. Notably, around the same time, Shanks argued that a value of $H_{0}\lesssim 50$ might permit a simpler inflationary model with $\Omega_{\text{baryon}}=1$, i.e. without invoking dark energy or cold dark matter [99].
@@
Notably, strong degeneracies in the parameter space related to the CMB have been reported recently. In [100] Alestas et al found that the best-fit value of $H_{0}$ obtained from the CMB power spectrum is degenerate with a constant equation of state (EoS) parameter $w$; the relationship is approximately linear, given by $H_{0}+30.93\,w-36.47=0$ (with $H_{0}$ in km sec-1 Mpc-1). Although this finding is not directly related to the BDRS work, the $H_{0}$-vs-$w$ degeneracy reinforces the general conclusion regarding the sensitivity of $H_{0}$ to other exogenous underlying assumptions that supplement the $\Lambda$CDM model—in the case of Alestas et al, the EoS parameter $w$.
@@
However, we believe that BDRS’s 2006 SDSS analysis contained an oversight, in light of our VSL cosmology. The standard Lemaître redshift formula and the conventional Hubble law are not applicable in the presence of varying speed of light. As discussed in Sections V.1 and V.2, these expressions are modified by a factor of $3/2$ due to varying $c$. Therefore, the SDSS would need a reevaluation to incorporate this VSL-induced adjustment. Here, we tentatively make a rudimentary fix:  we correct the comoving distance $s$ (measured in multiples of $h$, defined as $H_{0}/(100$ km sec-1 Mpc-1)), downward by a factor of $3/2$.In the lower panel of Fig. 8, we adjust the SDSS data (and their error bars) by reducing $s$ by a factor of $3/2$ and $s^{2}\,\xi$ by a factor of $(3/2)^{2}$. Upon these adjustments, the peak (at $s\simeq 75$) and trough (at $s\simeq 60$) of the SDSS become aligned with those of the BDRS’s model (dashed-dotted line), thereby lessening the discrepancy issue that led to BDRS’s abandonment of their original proposal.
@@ X Toward a new resolution of the  H0 tension
Formulae (70) and (71) thus allow for a current-time $H_{0}(z)$ ‘running’ as a function of the redshift of the data that are used to estimate it. With the function $F(z)$ parametrized in Eq. (61) with $b_{1}=b_{2}=2$ and $F_{\infty}=0.931$ as produced in Section VI.4, $H_{0}(z)$ can be computed using Eq. (71), as displayed in Fig. 9. At first, $H_{0}(z)$ decreases from $47.2$ (at $z=0$) to $41.5$ (at $z\simeq 2$), experiencing a 12% reduction. For $z\gtrsim 2$, $H_{0}(z)$ slowly rerises. At $z\rightarrow\infty$, with $F_{\infty}=0.931$ and $H_{0}(z=0)=47.2$, per Eq. (71), $H_{0}(z)$ asymptotically approaches $F_{\infty}H_{0}=43.95$, representing an 7% reduction from $H_{0}(z=0)$.
@@
    Remarkably, the asymptotic value $H_{0}(z\rightarrow\infty)=43.95$ that we just derived agrees surprisingly well with the $H_{0}\approx 44$ value obtained by Hunt and Sarkar in their follow-up study of the CMB [5].
@@ XI Discussions and summary
  • The Hubble constant $H_{0}=47.2\pm 0.4$ (95% CL). This value of consistent with the $3/2$ reduction referenced in Point (iii) above.
@@
III) Resolving the age problem. The age of an EdS universe is given by: $t_{0}=2/(3H_{0})$. Using the reduced value of $H_{0}=47.22\pm 0.4$ km sec-1 Mpc-1, one obtains $t_{0}=13.82\pm 0.11$ billion years. The age problem is thus resolved through the reduction in $H_{0}$, without requiring a recent acceleration phase induced by dark energy. See Section IX.
@@ Acknowledgements.
I thank Clifford Burgess, Tiberiu Harko, Robert Mann, Anne-Christine Davis and Eoin Ó Colgáin for their constructive and supportive comments during the development of this work.

### v5

@@ II A new mechanism to generate varying c from dilaton dynamics
Einstein’s pioneering concept of VSL, nevertheless, was quickly overshadowed by the success of his GR and subsequently fell into dormancy for several decades. The variability of $c$ was briefly rediscovered by Dicke in 1957 [11], prior to his own development of Brans–Dicke gravity [12], which instead allowed Newton’s gravitational constant $G$ to vary. In the 1990s, the idea of VSL was independently revived by Moffat [13] and by Albrecht and Magueijo [14] in the context of early-time cosmology. Their proposals aimed to resolve the horizon puzzle while avoiding the need for cosmic inflation. Since then, several researchers actively explore various aspects of VSL [33, 34, 35, 17, 27, 30, 29, 28, 18, 19, 49, 31, 32, 47, 48, 65, 80, 68, 45, 44, 43, 63, 62, 37, 36, 58, 72, 70, 75, 51, 71, 46, 25, 76, 81, 77, 73, 78, 55, 64, 66, 69, 39, 38, 52, 74, 79, 40, 53, 50, 54, 82, 26, 24, 22, 67, 61, 60, 21, 20, 16, 15, 57, 56, 23, 42, 41, 59, 83, 84, 85, 86, 87, 88].
@@ III.2 The modified FLRW metric
Here, $a_{0}$ is the current cosmic scale factor (often set equal 1), and $c_{0}$ is the speed of light measured at our current time in the intergalactic space. We should emphasize that the value of $c_{0}$ is not identical with the one measured inside the Milky Way, which is equal to $300,000$ km/s. This is because the Milky Way is a gravitationally bound structure whereas the intergalactic space is regions subject to cosmic expansion. This issue will be explained in Section IV.
@@ V.4 Modifying the luminosity distance–redshift formula
$4\pi r^{2}J=L\frac{\hat{\lambda}_{em}}{\hat{\lambda}_{ob}}.\frac{\hat{\lambda}_{em}}{\hat{\lambda}_{ob}}$ (47)
In the right hand side of Eq. (47), the first term $\hat{\lambda}_{em}/\hat{\lambda}_{ob}$ represents the “loss” in energy of the redshifted photon known as the “Doppler theft”  [^2] Note: This energy loss is consistent with the scaling property of energy for our VSL mechanism, as described by Eq. (13) in Section II. In intergalactic space, the energy of the traveling photon scales as $E\propto\chi\propto a^{-1}$, leading to a decline in energy as the universe expands.. The second (identical) term $\hat{\lambda}_{em}/\hat{\lambda}_{ob}$ arises from the dilution factor in photon density, as the same number of photons is distributed over a prolonged wavecrest in the radial direction (i.e., along the light ray). The $4\pi r^{2}$ in the left hand side of Eq. (47) accounts for the spherical dilution in flat space. From (46) and (47), we obtain
$d_{L}=r\,\frac{\hat{\lambda}_{ob}}{\hat{\lambda}_{em}}$ (48)
@@
where $\hat{d}_{L}$ is the luminosity distance observed by the Earth-based astronomer and $\hat{c}_{ob}$ the speed of light measured in the Milky Way (i.e., $300,000$ km/s). Formula (52) contains a single parameters $H_{0}$ and involves a function $F(z)$ that captures the evolution of the local scale of galaxies as a function of redshift.
@@ VI.1 $\Lambda$CDM and standard EdS as benchmarking models
with the sum taken over all $N=1,048$ Pantheon data points. The best fit for the $\Lambda$CDM model yields $H_{0}=70.2$ km/s/Mpc, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$, with the minimum error $\chi_{\text{min}}^{2}(\Lambda\text{CDM})=0.98824$. The $d_{L}$$z$ curve for the $\Lambda$CDM model is depicted by the dashed line in Fig. 3.
@@ VI.2 Fitting with VSL model: Disabling $F(z)$
[…]
@@ VI.3 Enabling $F(z)$:   A binning approach
[…]
@@
VI.4 Enabling $F(z)$:   A parametrization approach
@@
The steady decline of $F^{(i)}$ across the 10 bins with respect to redshift suggests adopting the following parametrization for $F(z)$:
@@
[…]
@@ VI.5 The cause for the reduction in $H_{0}$ value
Using this new function, BDRS produced an excellent fit to the CMB power spectrum, resulting in the following parameters: $H_{0}=46$ km/s/Mpc, $\omega_{\text{baryon}}:=\Omega_{\text{baryon}}(H_{0}/100)^{2}=0.019$, $\tau=0.16$ (the optical depth to last scattering), $k_{*}=0.0096$ Mpc-1, $n_{1}=1.015$, and $n_{2}=0.806$. The most remarkable outcome of the BDRS work is the “low” value of $H_{0}=46$, representing a 34% reduction from the accepted value of $H_{0}\sim 70$. A detailed follow-up study by Hunt and Sarkar [5], based on a supergravity-induced multiple inflation scenario, yielded a comparable value of $H_{0}\approx 44$. Notably, around the same time, Shanks argued that a value of $H_{0}\lesssim 50$ might permit a simpler inflationary model with $\Omega_{\text{baryon}}=1$, i.e. without invoking dark energy or cold dark matter [101].
@@
Notably, strong degeneracies in the parameter space related to the CMB have been reported recently. In [102] Alestas et al found that the best-fit value of $H_{0}$ obtained from the CMB power spectrum is degenerate with a constant equation of state (EoS) parameter $w$; the relationship is approximately linear, given by $H_{0}+30.93\,w-36.47=0$ (with $H_{0}$ in km/s/Mpc). Although this finding is not directly related to the BDRS work, the $H_{0}$-vs-$w$ degeneracy reinforces the general conclusion regarding the sensitivity of $H_{0}$ to other exogenous underlying assumptions that supplement the $\Lambda$CDM model—in the case of Alestas et al, the EoS parameter $w$.
@@
However, we believe that BDRS’s 2006 SDSS analysis contained an oversight, in light of our VSL cosmology. The standard Lemaître redshift formula and the conventional Hubble law are not applicable in the presence of varying speed of light. As discussed in Sections V.1 and V.2, these expressions are modified by a factor of $3/2$ due to varying $c$. Therefore, the SDSS would need a reevaluation to incorporate this VSL-induced adjustment. Here, we tentatively make a rudimentary fix:  we correct the comoving distance $s$ (measured in multiples of $h$, defined as $H_{0}/(100$ km/s/Mpc)), downward by a factor of $3/2$.In the lower panel of Fig. 8, we adjust the SDSS data (and their error bars) by reducing $s$ by a factor of $3/2$ and $s^{2}\,\xi$ by a factor of $(3/2)^{2}$. Upon these adjustments, the peak (at $s\simeq 75$) and trough (at $s\simeq 60$) of the SDSS become aligned with those of the BDRS’s model (dashed-dotted line), thereby lessening the discrepancy issue that led to BDRS’s abandonment of their original proposal.
@@ X Toward a new resolution of the  H0 tension
Formulae (70) and (71) thus allow for a current-time $H_{0}(z)$ ‘running’ as a function of the redshift of the data that are used to estimate it. With the function $F(z)$ parametrized in Eq. (61) with $b_{1}=b_{2}=2$ and $F_{\infty}=0.931$ as produced in Section VI.4, $H_{0}(z)$ can be computed using Eq. (71), as displayed in Fig. 9. At first, $H_{0}(z)$ decreases from $47.2$ (at $z=0$) to $41.5$ (at $z\simeq 2$), experiencing a 12% reduction. For $z\gtrsim 2$, $H_{0}(z)$ slowly rerises. At $z\rightarrow\infty$, with $F_{\infty}=0.931$ and $H_{0}(z=0)=47.22$, per Eq. (71), $H_{0}(z)$ asymptotically approaches $F_{\infty}H_{0}=43.96$, representing an 7% reduction from $H_{0}(z=0)$.
@@
    Remarkably, the asymptotic value $H_{0}(z\rightarrow\infty)=43.96$ that we just derived agrees surprisingly well with the $H_{0}\approx 44$ value obtained by Hunt and Sarkar in their follow-up study of the CMB [5].
@@ XI Discussions and summary
  • The Hubble constant $H_{0}=47.22\pm 0.4$ (95% CL). This value of consistent with the $3/2$ reduction referenced in Point (iii) above.
@@
Our modified Lemaître redshift formula, Eq. (35), can also effectively viewed as a form of “redshift remapping”, a technique advocated in Refs. [112, 113, 111]. Interestingly, our value of $H_{0}=47.2\pm 0.4$ aligns with the result $H_{0}=48\pm 2$ reported in [113].
@@
III) Resolving the age problem. The age of an EdS universe is given by: $t_{0}=2/(3H_{0})$. Using the reduced value of $H_{0}=47.22\pm 0.4$ km/s/Mpc, one obtains $t_{0}=13.82\pm 0.11$ billion years. The age problem is thus resolved through the reduction in $H_{0}$, without requiring a recent acceleration phase induced by dark energy. See Section IX.
@@
On the cosmological time dilation extracted from the Dark Energy Survey (DES)
A recent paper [114] using DES supernova light curves showed no deviation from the relation $\Delta t_{obs}=\Delta t_{em}(1+z)$. However, this finding does not contradict our modified Lemaitre redshift formula, $1+z=a^{-3/2}F(z)$. This is because the result in Ref. [114] only verifies that the speed of light inside the galaxies hosting the supernovae and that inside the Milky Way are approximately the same. Galaxies are gravitationally bound and thus not subject to cosmic expansion. Reference [114] does not deal with the speed of light in the intergalactic space, the expansion of which causes c to decline over time. We clarified this distinction in Section IV.
XII Conclusion
@@
Finally, we note that the observational bounds established in the literature in support of a constant speed of light have predominantly relied on standard cosmology [41, 42, 60, 61, 82, 34, 35, 36, 37, 43, 44, 45, 68, 33, 67, 65, 71]. However, our new Lemaître redshift formula represents a critical departure from this conventional framework. Therefore, the consensus regarding the absence of variation in $c$ in observational cosmology must be reconsidered in light of our findings, prompting a comprehensive reanalysis of these constraints.
@@ Acknowledgements.
I thank Clifford Burgess, Tiberiu Harko, Robert Mann, Anne-Christine Davis, Eoin Ó Colgáin, Demosthenes Kazanas, Leandros Perivolaropoulos, John Moffat, Alberto Salvio, Ilya Shapiro, Gregory Volovik and Bruce Bassett for their constructive feedback.